# A refined Weyl character formula for comodules on $\operatorname{GL}_{2,A}$

### v4

@@ 1. Introduction
The quotient in 1.1 is proved to exist in [8], [3] and [13] over an arbitrary Dedekind domain, but the ring $A$ in Theorem 1.1 may be an arbitrary commutative unital ring. I give an explicit construction of the representing $A$-Hopf algebra $k[N/T]$ of $N/T$ and the morphism $\pi$.
@@
In Example 6.6 we relate the refined weight space decomposition to irreducible finite dimensional modules over a field $k$ of positive charcacteristic. If $G:=\operatorname{SL}_{n,k}$ and if $S_{\lambda}(V)$ is a Weyl module, there is a unique finite dimensional irreducible module $V(\lambda)\subsetneq S_{\lambda}(V)$. The irreducible module $V(\lambda)$ has a relation with the refined weight space decomposition for $S_{\lambda}(V)$.
In Theorem 7.1 I prove that the direct summand $\operatorname{Sym}^{d}(V)^{i}$ is an irreducible module on $N$ over any field $k$.
@@
It is hoped the results in this paper may be interesting in the study of torsion free comodules on $\operatorname{SL}_{2,A}$ and $\operatorname{GL}_{2,A}$ in the future. One problem one wants to study is the problem of classifying $\operatorname{SL}_{2,\mathbf{Z}}$ modules $W$ whose base change to $\operatorname{SL}_{2,\mathbf{Q}}$ are isomorphic to $\operatorname{Sym}^{d}(\mathbf{Q}^{2})$ where $\mathbf{Q}^{2}$ is the standard module on $\operatorname{SL}_{2,\mathbf{Q}}$. Such a classification may have interest in the problem of giving a definition of the notion ”good filtration” for torsion free finite rank comodules on $\operatorname{SL}_{2,\mathbf{Z}}$. Such a notion may have applications in the study of the representation theory of $\operatorname{SL}_{n,k}$ and $\operatorname{GL}_{n,k}$ over an arbitrary field $k$.
@@ 5. On the schematic normalizer in $\operatorname{GL}_{2,A}$
where the right hand side is the semi direct product with respect to $\sigma_{B}$. The map $\sigma_{B}$ is functorial and we get a natural transformation $\sigma:h_{N}(-)\rightarrow h_{N/T}(-)$ of functors.
@@
A relation with irreducible modules over a field of characteristic $p>0$.
@@
and the irreducible module $V(2)$ with highest weight $2$ is the strict submodule $V(2):=k\{e_{1}^{2},e_{2}^{2}\}\subsetneq\operatorname{Sym}^{2}(V)$. Hence there is an equality
@@
where $\operatorname{Sym}^{2}(V)^{1}$ is a term in the refined weight space decomposition. Hence the refined weight space decomposition contains information on irreducible modules in positive characteristic.
@@ 7. On irreducibility and the refined weight space decomposition
Irreducible finite dimensional modules over a field of characteristic $p>0$.
If $G:=\operatorname{GL}(V)$ or $\operatorname{SL}(V)$ for $V:=k^{n}$ an $n$-dimensional vector space over $k$, it follows for any finite dimensional irreducible $G$-module $V(\lambda)$, there is a Weyl module $S_{\lambda}(V)$ and an inclusion $V(\lambda)\subseteq S_{\lambda}(V)$ which is strict in general. It is an open problem to calculate the character $Char(V(\lambda))$. If the terms in the refined weight space decomposition are irreducible modules over the normalizer $N$, this decomposition may be interesting in the study of such character formulas. There are canonical diract summand decompositions
@@
and if $S_{\lambda}(V)^{j}$ is an irreducible module, it follows the inclusion map

### v5

@@ 1. Introduction
The quotient in 1.1 is proved to exist in [8], [3] and [13] over an arbitrary Dedekind domain, but the ring $A$ in Theorem 1.1 may be an arbitrary commutative unital ring. I give an explicit construction of the representing $A$-Hopf algebra $k[N/T]$ of $N/T$ and the morphism $\pi$. If $W_{G}$ is the Weyl group of $\operatorname{GL}_{2,A}$ it follows there is an isomorphism of group schemes
$N\cong T\times^{\sigma}\operatorname{Spec}(k[W_{G}])$
where $\operatorname{Spec}(k[W_{G}])$ is the constant group scheme of the Weyl group $W_{G}$.
@@
In Example 6.6 we relate the refined weight space decomposition to irreducible finite dimensional comodules over a field $k$ of positive charcacteristic. If $G:=\operatorname{SL}_{n,k}$ and if $S_{\lambda}(V)$ is a Weyl module, there is a unique finite dimensional irreducible comodule $V(\lambda)\subsetneq S_{\lambda}(V)$. The irreducible comodule $V(\lambda)$ has a relation with the refined weight space decomposition for $S_{\lambda}(V)$.
In Theorem 7.1 I prove that the direct summand $\operatorname{Sym}^{d}(V)^{i}$ is an irreducible comodule on $N$ over any field $k$. In general let $V$ be a finite dimensional vector space over a field $k$ of characteristic $p>0$ and let $S_{\lambda}(V)$ be the Weyl module of $V$ corresponding to the partition $\lambda$. If $G:=\operatorname{GL}(V)$ there is in general a unique irreducible sub-$G$-comodule
$V(\lambda)\subseteq S_{\lambda}(V).$
There are refined weight space decompositions $S_{\lambda}(V)\cong\oplus_{i\in I}S_{\lambda}(V)^{i}$ and $V(\lambda)\cong\oplus_{j\in J}V(\lambda)^{j}$ and there are relations between these two decompositions. In the case when $V\cong k^{2}$ we may express the irreducible comodule $V(\lambda)$ in terms of the refined weight space decomposition for $\operatorname{Sym}^{d}(V)$. Hence the construction may have applications in the study of irreducible $G$-comodules over any field $k$.
@@
It is hoped the results in this paper may be interesting in the study of torsion free comodules on $\operatorname{SL}_{2,A}$ and $\operatorname{GL}_{2,A}$ in the future. One problem one wants to study is the problem of classifying $\operatorname{SL}_{2,\mathbf{Z}}$ comodules $W$ whose base change to $\operatorname{SL}_{2,\mathbf{Q}}$ are isomorphic to $\operatorname{Sym}^{d}(\mathbf{Q}^{2})$ where $\mathbf{Q}^{2}$ is the standard comodule on $\operatorname{SL}_{2,\mathbf{Q}}$. Such a classification may have interest in the problem of giving a definition of the notion ”good filtration” for torsion free finite rank comodules on $\operatorname{SL}_{2,\mathbf{Z}}$. Such a notion may have applications in the study of the representation theory of $\operatorname{SL}_{n,k}$ and $\operatorname{GL}_{n,k}$ over an arbitrary field $k$.
@@ 5. On the schematic normalizer in $\operatorname{GL}_{2,A}$
where the right hand side is the semi direct product with respect to $\sigma_{B}$. The map $\sigma_{B}$ is functorial and we get a natural transformation $\sigma:h_{N/T}(-)\rightarrow h_{N}(-)$ of functors.
@@
A relation with irreducible comodules over a field of characteristic $p>0$.
@@
and the irreducible comodule $V(2)$ with highest weight $2$ is the strict sub-comodule $V(2):=k\{e_{1}^{2},e_{2}^{2}\}\subsetneq\operatorname{Sym}^{2}(V)$. Hence there is an equality
@@
where $\operatorname{Sym}^{2}(V)^{1}$ is a term in the refined weight space decomposition. Hence the refined weight space decomposition contains information on irreducible comodules in positive characteristic.
@@ 7. On irreducibility and the refined weight space decomposition
Irreducible finite dimensional comodules over a field of characteristic $p>0$.
If $G:=\operatorname{GL}(V)$ or $\operatorname{SL}(V)$ for $V:=k^{n}$ an $n$-dimensional vector space over $k$, it follows for any finite dimensional irreducible $G$-comodule $V(\lambda)$, there is a Weyl module $S_{\lambda}(V)$ and an inclusion $V(\lambda)\subseteq S_{\lambda}(V)$ which is strict in general. It is an open problem to calculate the character $Char(V(\lambda))$. If the terms in the refined weight space decomposition are irreducible comodules over the normalizer $N$, this decomposition may be interesting in the study of such character formulas. There are canonical diract summand decompositions
@@
and if $S_{\lambda}(V)^{j}$ is an irreducible comodule, it follows the inclusion map